OpenStudy (ujjwal):

How does a PNP transistor work? And what does it do? I need few concepts regarding it. I have already studied junction diodes..

OpenStudy (shayaan_mustafa):

Hi ujjwal :) May I help you as it belongs to my field. Electronics engineering.

OpenStudy (ujjwal):

sure

OpenStudy (ujjwal):

Of course i need help.. you must be an expert in such topics.

OpenStudy (shayaan_mustafa):

So you have studied diode So you know transistors are two opposite diodes.

OpenStudy (ujjwal):

yeah. NPN and PNP.. Combination of N and P type semiconductor.

OpenStudy (anonymous):

its mainly amplification device.... here's too much wid transistor buddy, r u want full explaination?

OpenStudy (ujjwal):

I just need the basic ideas.. I will be glad if in the end i am able to differentiate and calculate base and emitter current.

OpenStudy (shayaan_mustafa):

Have diagram of transistor? Or just keep it in mind. you know we require beta i.e. current gain in order to calculate current of base, collector or emitter?

OpenStudy (ujjwal):

 can you explain what beta is?

OpenStudy (shayaan_mustafa):

beta is the current. It vary from transistor to transistor. It is pre-defined on the specification sheet or you say data sheet. \[\large I_\beta=\frac{I_C}{I_B}\] Where, \[I_\beta=\text{Current gain}\]\[I_C=\text{Collector current}\]\[I_B=\text{Base current}\] This formula is same for PNP and NPN transistor.

OpenStudy (shayaan_mustafa):

Does this help you dear?

OpenStudy (ujjwal):

wait a min. I have a question.

OpenStudy (shayaan_mustafa):

Yes yes sure.

OpenStudy (ujjwal):

The value of \(\alpha\)=0.9, the change in collector current in transistor if base current changes by 2 micro Amp is:

OpenStudy (shayaan_mustafa):

Oh so you are confusing between alpha and beta. right?

OpenStudy (anonymous):

\[\alpha = I _c{}/ I _e {}\]

OpenStudy (shayaan_mustafa):

Alpha is actually known as common base current gain Beta is actually known as common emitter current gain But both are current gain. I have showed you beta above. Here is alpha \[\large \alpha=\frac{I_C}{I_E}\] where, \[I_C=\text{collector current}\]\[\large I_E=\text{emitter current}\]

OpenStudy (anonymous):

\[\beta = I _c{}/ I _b{}\]

OpenStudy (shayaan_mustafa):

Yes kr7210's equations are right.

OpenStudy (anonymous):

\[\beta = \alpha /1-\alpha \]

OpenStudy (anonymous):

u can get ur answer by using all three formulas

OpenStudy (ujjwal):

Is there a theoretical interpretation of those formulas? I can mug them up but i don't know the theory part. still unaware of basic ideas..

OpenStudy (anonymous):

too difficult to explain here buddy... its take too much tym to write somthng

OpenStudy (ujjwal):

ok. anyways thanks! i will come up with specific problems soon..

OpenStudy (shayaan_mustafa):

@ujjwal still need help dear?